4x + 2y =60 solve for x
Answer by ptaylor(2198)

4x + 2y =60 solve for x

First, subtract 2y from both sides
4x+2y-2y=60-2y combine like terms
4x=60-2y next, divide both sides by 4
x=(60-2y)/4 divide numerator and denominator by 2
x=(30-y)/2-----------ans

Hope this helps----ptaylor